# Do Language Models Associate Sound with Meaning? A Multimodal Study of Sound Symbolism

### v4

@@ A.2 Semantic Dimension
Table 6 shows the extreme feature pairs of 25 semantic dimensions suggested by sidhu2022higher, which have been utilized for our semantic dimension prediction experiments.
@@
Table 6: The 25 semantic dimension list by sidhu2022higher.
@@ B.1 Inference Settings
  • Software: PyTorch (paszke2017automatic) and vLLM (kwon2023efficient) libraries.
@@ B.2 Language Models Used in This Work
  • gpt-4o (openai2024gpt4ocard)
  • gpt-4o-audio-preview (openai2024gpt4ocard)
  • gpt-4.1 (openai2025_gpt41)
  • gemini-2.5-flash (comanici2025gemini25pushingfrontier)
  • gemma-3-27b-it (gemmateam2025gemma3technicalreport)
  • Qwen3-32B (yang2025qwen3technicalreport)
  • Qwen2.5-Omni-3B (xu2025qwen25omnitechnicalreport)
  • Qwen2.5-Omni-7B (xu2025qwen25omnitechnicalreport)
@@ C.1 Montreal Forced Alignment (MFA) Pipeline
We employ Montreal Forced Alignment (mcauliffe2025montreal) to obtain precise phoneme-level segmentation of audio waveforms. We utilize pre-trained acoustic models for each language:
  • English (US) ARPA acoustic model v3.0.0 (mfa_english_us_arpa_acoustic_2024)
  • French MFA acoustic model v3.0.0 (mfa_french_mfa_acoustic_2024)
  • Japanese MFA acoustic model v3.0.0 (mfa_japanese_mfa_acoustic_2024)
  • Korean MFA acoustic model v3.0.0 (mfa_korean_mfa_acoustic_2024)
@@ Appendix E Human Evaluation
To ensure the reliability of the semantic dimension selection and the automatic annotation of the pseudo ground truth data in LEX-ICON, we conduct a human evaluation using 152 questions randomly sampled for each participant from the entire dataset. 10 graduate students without expertise in linguistics participate in this experiment, which is performed via the Label Studio (LabelStudio). The experiment utilize questions with the same prompt structure as the semantic dimension prediction test that was provided to LLMs. With regard to the experiment's feasibility, all words are presented in audio waveforms directly hearable to humans. An example image of the process is displayed in Figure 10.

### v5

@@ A.2 Semantic Dimension
Table 6 shows the extreme feature pairs of 25 semantic dimensions suggested by Sidhu et al. (2022), which have been utilized for our semantic dimension prediction experiments.
@@
Table 6: The 25 semantic dimension list by Sidhu et al. (2022).
@@ B.1 Inference Settings
  • Software: PyTorch (Paszke et al. 2017) and vLLM (Kwon et al. 2023) libraries.
@@ B.2 Language Models Used in This Work
  • gpt-4o (OpenAI et al. 2024)
  • gpt-4o-audio-preview (OpenAI et al. 2024)
  • gpt-4.1 (OpenAI 2025)
  • gemini-2.5-flash (Comanici et al. 2025)
  • gemma-3-27b-it (Team et al. 2025b)
  • Qwen3-32B (Yang et al. 2025a)
  • Qwen2.5-Omni-3B (Xu et al. 2025)
  • Qwen2.5-Omni-7B (Xu et al. 2025)
@@ C.1 Montreal Forced Alignment (MFA) Pipeline
We employ Montreal Forced Alignment (McAuliffe and others 2025) to obtain precise phoneme-level segmentation of audio waveforms. We utilize pre-trained acoustic models for each language:
  • English (US) ARPA acoustic model v3.0.0 (McAuliffe and Sonderegger 2024a)
  • French MFA acoustic model v3.0.0 (McAuliffe and Sonderegger 2024b)
  • Japanese MFA acoustic model v3.0.0 (McAuliffe and Sonderegger 2024c)
  • Korean MFA acoustic model v3.0.0 (McAuliffe and Sonderegger 2024d)
@@ Appendix E Human Evaluation
To ensure the reliability of the semantic dimension selection and the automatic annotation of the pseudo ground truth data in LEX-ICON, we conduct a human evaluation using 152 questions randomly sampled for each participant from the entire dataset. 10 graduate students without expertise in linguistics participate in this experiment, which is performed via the Label Studio (Tkachenko et al. 2020). The experiment utilize questions with the same prompt structure as the semantic dimension prediction test that was provided to LLMs. With regard to the experiment's feasibility, all words are presented in audio waveforms directly hearable to humans. An example image of the process is displayed in Figure 10.